# Systolic Arrays and Structured Pruning Co-design for Efficient Transformers in Edge Systems

### v1

@@ IV-B Hardware Exploration
Figure 6: Synthesis results for the systolic array design across configurations of varying size (number of rows of the array) and quantization.
@@
IV-C System Exploration
Figure 7: Per-layer normalized run-time of the transformer encoder after applying Systolic Array Structured Pruning at two global sparsity targets. Results are shown for an $8\times 8$, FP32_INT8 systolic array, with the execution time of each layer normalized to the execution without pruning.
Figure 7 plots the measured per-layer normalized encoder run-time of a systolic-accelerated system performing an inference. Data refers to an $8\times 8$ systolic array, with varying degrees of structured sparsity. Speedup numbers closely follow sparsity levels, as inference run-time is strongly dominated by GEMM computations (exceeding 97% in all cases [20]).
Pruning was performed according to the methodology proposed in Section III-A targeting feed-forward layers, which exhibit a much higher degree of resilience and account for the largest part of the workload. Results in Figure 7 highlight that early feed-forward layers are the most amenable to pruning, while later ones have a higher proportion of tiles with a non-negligible L1-norm, which have a higher impact on inference outcomes.
IV-D Impact of SASP on Quality of Service
@@
Figure 8: Achieved Word Error Rate when varying the percentage of Systolic Array Structured Pruning.
@@
IV-E Multidimensional SASP Trade-offs
Figure 9: Trade-offs among encoder inference speedup, area-energy product and Word Error Rate across systolic array sizes and structured pruning rates. Speedup is computed with respect to a non-accelerated, non-quantized baseline.
Figure 9 shows the variations in performance, WER and resource usage when changing SASP rate, quantization strategy, and systolic array size. WER and Speedup are plotted on the two axes, the marker shape discriminates between FP32 and weight-quantized implementations, and the marker colors indicate their resource requirements in terms of Area-Energy product. Data points form four distinct clusters corresponding to each systolic array size. Two curves in each cluster represent the two quantization choices, which have notably different Area-Energy products. The non-quantized FP32_FP32 version achieves lower (better) WER, with the differences becoming more pronounced at higher pruning rates and for larger systolic arrays.
@@ IV-E Multidimensional SASP Trade-offs
Within each systolic array size and quantization configuration, the SASP pruning rate guides the trade-off between inference time and WER. Figure 9 shows that, up to an inflection point at a WER of $\sim$5%, SASP enables to strike effective fine-grained balances between run-time performance and QoS. Beyond this inflection point, further increases in pruning rates cause instead very high WER degradations.
Table III illustrates the effect of applying Systolic Array Structured Pruning and weight quantization at the 5% inflection point. In this setting, SASP improves performance and energy consumption up to 26% and 21%, respectively. Furthermore, when combining quantization and structured pruning, performance and energy efficiency improvements reach 44% and 42%, while also decreasing area occupation by 36%. These substantial gains are achieved without increasing the systolic array size, and hence do not incur the hefty associated area and energy costs, as also depicted in Table III. As an example, scaling from an $8\times 8$ to a $32\times 32$ systolic array does yield a 3.04$\times$ speedup for FP32_INT8 quantization, but also requires 15.21$\times$ more area and 3.98$\times$ more energy.
TABLE III: Area, encoder speedup and energy results for different systolic array configurations without SASP (3.5$\pm$0.2WER) and with SASP (5$\pm$0.4WER). Speedup is computed with respect to a non-quantized baseline executed on CPU.
@@
IV-F Cross-tier Analysis
As discussed above, increasing the systolic array size increases run-time performance by offering higher parallelism, but also restricts the achievable pruning rate for a given WER. Therefore, when using SASP, increases in the array size (hence the tile size in GEMM computations) result in diminishing gains for a given WER target, as the pruning rate must be lowered to maintain QoS. This trend is illustrated in Figure 10, which highlights a sublinear relation between systolic array size and speedup, both in the case of FP32 and weight-quantized scenarios. Hardware costs (area / energy), instead, increase quadratically with size, as outlined in Section IV-B, requiring careful co-design considerations, especially when targeting resource-constrained edge systems.
@@ IV-F Cross-tier Analysis
Figure 10: Speedup with respect to software execution of the encoder while varying the systolic array size, for different Word Error Rates.
V Conclusion
Deploying transformers on edge devices requires both software optimization and hardware acceleration to meet strict resource constraints while maintaining performance. In this paper, we have analyzed their interaction, focusing on structured sparsity and systolic arrays. We explored Systolic Array Structured Pruning (SASP), where the size of pruned blocks is matched to the dimensions of the systolic array, enabling the skipping of entire computation tiles. To assess the benefits and pitfalls of SASP, we presented a cross-stack framework to co-optimize edge AI transformers, which integrates algorithmic optimization, system simulation, and hardware design. Employing it, we performed a comprehensive analysis of how SASP, quantization, and systolic array configurations affect area, energy, performance, and Word Error Rate (WER) in edge transformers for speech recognition.
Our results demonstrate that SASP provides fine-grained control over the trade-off between inference run-time and WER up to an inflection point, after which increased pruning drastically degrades QoS for small performance gains. Experimental evidence has shown that system-wide performance improvements of up to 44% and accelerator energy reductions of up to 42% can be obtained under a 1.4% WER degradation, when employing weight quantization and a 20% pruning rate. Additionally, we showcased that, although larger systolic arrays do reduce run-time, they also incur substantial energy and area costs, while yielding sublinear speedups for a target WER. This sublinearity emerges from the reduced structured pruning opportunities, as finding contiguous zero blocks becomes harder with increasing block sizes. For this reason, SASP is particularly well-suited for edge AI accelerators, where stringent resource constraints are present.

### v2

@@ 4.2. Hardware Exploration
Figure 6. Synthesis results for the systolic array design across configurations of varying size (number of rows of the array) and quantization.
@@
4.3. System Exploration
Figure 7. Speedup and energy improvements from Systolic Array Structured Pruning under the quality of service target in Table 1 with respect to non-pruned quantized executions, considering systems with systolic array accelerators of varying size. Results are shown for FP32_INT8 systolic array configurations.
We first assess the overall impact of SASP across workloads and systolic array sizes on runtime and energy consumption under the target QoS degradations defined in Table 1. As shown in Figure 7, speedup and energy improvement opportunities depend on the workload due to their different sensitivities to pruning. For example, maximum speedup / energy improvements vary from 26% / 21% in ESPnet ASR on LibriSpeech, to 22% / 18% in ESPnet2 ASR on LibriSpeech, to 51% / 34% in ESPnet2 ASR+MT on MuST-C. However, their trends when changing accelerator size are consistent across workloads, e.g., achievable improvements decrease when employing larger systolic arrays. To thoroughly analyze the trade-offs among encoder inference speedup, area-energy product and quality of service, all further explorations will focus on the ESPnet automatic speech recognition model using the LibriSpeech corpus.
Figure 8 plots the measured per-layer normalized encoder run-time of a systolic-accelerated system performing an inference. Data refers to an $8\times 8$ systolic array, with varying degrees of structured sparsity. Speedup numbers closely follow sparsity levels, as inference run-time is strongly dominated by GEMM computations (exceeding 97% in all cases (Amirshahi et al., 2023)).
Figure 8. Per-layer normalized run-time of the transformer encoder after applying Systolic Array Structured Pruning at two global sparsity targets. Results are shown for an $8\times 8$, FP32_INT8 systolic array, with the execution time of each layer normalized to the execution without pruning.
Pruning was performed according to the methodology proposed in Section 3.1 targeting feed-forward layers, which exhibit a much higher degree of resilience and account for the largest part of the workload. Results in Figure 8 highlight that early feed-forward layers are the most amenable to pruning, while later ones have a higher proportion of tiles with a non-negligible L1-norm, which have a higher impact on inference outcomes.
4.4. Impact of SASP on Quality of Service
@@
Figure 9. Achieved Word Error Rate when varying the percentage of Systolic Array Structured Pruning.
@@
4.5. Multidimensional SASP Trade-offs
Figure 10. Trade-offs among encoder inference speedup, area-energy product and Word Error Rate across systolic array sizes and structured pruning rates. Speedup is computed with respect to a non-accelerated, non-quantized baseline.
Focusing again on speech recognition, Figure 10 shows the variations in performance, QoS and resource usage when changing SASP rate, quantization strategy, and systolic array size. WER and Speedup are plotted on the two axes, the marker shape discriminates between FP32 and weight-quantized implementations, and the marker colors indicate their resource requirements in terms of Area-Energy product. Data points form four distinct clusters corresponding to each systolic array size. Two curves in each cluster represent the two quantization choices, which have notably different Area-Energy products. The non-quantized FP32_FP32 version achieves lower (better) WER, with the differences becoming more pronounced at higher pruning rates and for larger systolic arrays.
@@ 4.5. Multidimensional SASP Trade-offs
Within each systolic array size and quantization configuration, the SASP pruning rate guides the trade-off between inference time and QoS. Figure 10 shows that, up to an inflection point at a WER of $\sim$5%, SASP enables to strike effective fine-grained balances between run-time performance and QoS. Beyond this inflection point, further increases in pruning rates cause instead very high WER degradations.
Table 3 illustrates the effect of applying Systolic Array Structured Pruning and weight quantization at the 5% inflection point. In this setting, SASP improves performance and energy consumption up to 26% and 21%, respectively. Furthermore, when combining quantization and structured pruning, performance and energy efficiency improvements reach 44% and 42%, while also decreasing area occupation by 36%. These substantial gains are achieved without increasing the systolic array size, and hence do not incur the hefty associated area and energy costs, as also depicted in Table 3. As an example, scaling from an $8\times 8$ to a $32\times 32$ systolic array does yield a 3.04$\times$ speedup for FP32_INT8 quantization, but also requires 15.21$\times$ more area and 3.98$\times$ more energy.
Table 3. Area, encoder speedup and energy results for different systolic array configurations without SASP (3.5$\pm$0.2WER) and with SASP (5$\pm$0.4WER). Speedup is computed with respect to a non-quantized baseline executed on CPU.
@@
4.6. Cross-tier Analysis
As discussed above, increasing the systolic array size increases run-time performance by offering higher parallelism, but also restricts the achievable pruning rate for a given QoS. Therefore, when using SASP, increases in the array size (hence the tile size in GEMM computations) result in diminishing gains for a given WER target, as the pruning rate must be lowered to maintain QoS. This trend is illustrated in Figure 11, which highlights a sublinear relation between systolic array size and speedup, both in the case of FP32 and weight-quantized scenarios. Hardware costs (area / energy), instead, increase quadratically with size, as outlined in Section 4.2, requiring careful co-design considerations, especially when targeting resource-constrained edge systems.
@@ 4.6. Cross-tier Analysis
Figure 11. Speedup with respect to software execution of the encoder while varying the systolic array size, for different Word Error Rates.
5. Conclusion
Deploying transformers on edge devices requires both software optimization and hardware acceleration to meet strict resource constraints while maintaining performance. In this paper, we have analyzed their interaction, focusing on structured sparsity and systolic arrays. We explored Systolic Array Structured Pruning (SASP), where the size of pruned blocks is matched to the dimensions of the systolic array, enabling the skipping of entire computation tiles. To assess the benefits and pitfalls of SASP, we presented a cross-stack framework to co-optimize edge AI transformers, which integrates algorithmic optimization, system simulation, and hardware design. Employing it, we performed a comprehensive analysis of how SASP, quantization, and systolic array configurations affect area, energy, performance, and quality of service (QoS) in edge transformers for speech recognition and machine translation.
Our results demonstrate that SASP provides fine-grained control over the trade-off between inference run-time and QoS up to an inflection point, after which increased pruning drastically degrades QoS for small performance gains. Experimental evidence on the ESPnet ASR transformer trained on LibriSpeech has shown that system-wide performance improvements of up to 44% and accelerator energy reductions of up to 42% can be obtained under a 1.4% WER degradation, when employing weight quantization and a 20% pruning rate. Additionally, we showcased that, although larger systolic arrays do reduce run-time, they also incur substantial energy and area costs, while yielding sublinear speedups for a target QoS. This sublinearity emerges from the reduced structured pruning opportunities, as finding contiguous zero blocks becomes harder with increasing block sizes. For this reason, SASP is particularly well-suited for edge AI accelerators, where stringent resource constraints are present.
Acknowledgements.
This work was supported in part by the Swiss State Secretariat for Education, Research, and Innovation (SERI) through the SwissChips Research Project; in part by the Swiss NSF Edge-Companions project (GA No. 10002812); in part by the EC H2020 FVLLMONTI Project under Grant 101016776; and in part by “La Caixa” Foundation (ID 100010434) under Grant LCF/BQ/EU23/12010073. This research was partially conducted by ACCESS – AI Chip Center for Emerging Smart Systems, supported by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government.